# Dual Refinement Cycle Learning: Unsupervised Text Classification of Mamba and Community Detection on Text Attributed Graph

### v1

@@ Abstract
Pretrained language models offer strong text understanding capabilities but remain difficult to deploy in real-world text-attributed networks due to their heavy dependence on labeled data. Meanwhile, community detection methods typically ignore textual semantics, limiting their usefulness in downstream applications such as content organization, recommendation, and risk monitoring. To overcome these limitations, we present Dual Refinement Cycle Learning (DRCL), a fully unsupervised framework designed for practical scenarios where no labels or category definitions are available.
DRCL integrates structural and semantic information through a warm-start initialization and a bidirectional refinement cycle between a GCN-based Community Detection Module (GCN-CDM) and a Text Semantic Modeling Module (TSMM). The two modules iteratively exchange pseudo-labels, allowing semantic cues to enhance structural clustering and structural patterns to guide text representation learning without manual supervision.
Across several text-attributed graph datasets, DRCL consistently improves the structural and semantic quality of discovered communities. Moreover, a Mamba-based classifier trained solely from DRCL’s community signals achieves accuracy comparable to supervised models, demonstrating its potential for deployment in large-scale systems where labeled data are scarce or costly.
@@
[a] organization=College of Physics and Information Engineering, Minnan Normal University, addressline=, city=Zhangzhou, state=Fujian, postcode=363000, country=China
@@ 3 Preliminaries
Definition 1. Text-Attributed Graphs. A text-attributed graph can be formally defined as a quadruple $G=(V,E,X,T)$, where $V=\{v_{1},v_{2},...,v_{n}\}$ denotes the set of nodes, and $n=|V|$ represents the total number of nodes in the graph. $E\subseteq V\times V$ is the set of edges, representing pairwise relationships between nodes. An edge $e_{ij}=(v_{i},v_{j})\in E$ indicates that there is a connection between node $v_{i}$ and node $v_{j}$. $M=|E|$ denotes the total number of edges. $X=\left\{x_{1},x_{2},x_{3},\ldots,x_{n}\right\}\in\mathbb{R}^{n\times d}$ is the node attribute feature matrix. The $i$-th row vector $x_{i}\in\mathbb{R}^{d}$ corresponds to the $d$-dimensional feature representation of node $v_{i}$. $T=\{t_{1},t_{2},...,t_{n}\}$ denotes the textual content associated with the $n$ nodes. The topology of the graph is described by the adjacency matrix $A\in\{0,1\}^{n\times n}$, where $a_{ij}=1$ if and only if $(v_{i},v_{j})\in E$; otherwise, $a_{ij}=0$. Based on the adjacency matrix, the degree matrix can be defined as $D=\operatorname{diag}\left(d_{1},d_{2},d_{3},\ldots,d_{n}\right)$. It is a diagonal matrix whose elements $d_{i}=\sum_{j=1}^{n}a_{ij}$ represent the degree of node $v_{i}$, that is, the number of neighbors directly connected to it. The degree matrix characterizes the structural information of the entire graph.
Definition 2. Graph Convolutional Network. The core idea of GCNs originates from spectral graph theory, where graph signals are filtered in the frequency domain. The layer-wise propagation rule is a first-order Chebyshev polynomial approximation of the graph Laplacian operator, enabling efficient neighborhood aggregation in the spatial domain. The propagation rule of GCN is given by:
@@
where $H^{(l)}=\{h_{1}^{(l)},h_{2}^{(l)},...,h_{n}^{(l)}\}\in\mathbb{R}^{n\times d`}$ is the node representation matrix at layer $l$, and the initial input is $H^{(0)}=X$. $\tilde{A}=\hat{D}^{-1/2}\hat{A}\hat{D}^{-1/2}$ is the normalized adjacency matrix, used to stabilize training and prevent gradient explosion or vanishing. $\hat{A}=A+I$ is the adjacency matrix with self-loops, where $I$ is the identity matrix, ensuring that each node preserves its own features during aggregation. $\hat{D}$ is the degree matrix of $\hat{A}$, defined as $\hat{D}_{ii}={\textstyle\sum_{j}}\hat{A}_{ij}$. $W^{(l)}$ is the learnable weight matrix at layer $l$. $\sigma(\cdot)$ is the PReLU activation function, which enhances nonlinear representation capacity and improves feature discrimination.
@@ 5.1 Datasets and Evaluation Metrics
To ensure the completeness and comparability of the evaluation, this paper selects six widely used metrics for comprehensive assessment: DBI, Q, NMI, ACC, F1, and ARI. Among these, DBI and Q measure the quality of community partitioning from the perspectives of internal clustering cohesion and graph structure consistency, respectively. NMI, ACC, F1, and ARI are external consistency metrics commonly used in both community detection and language text classification tasks, reflecting the degree of alignment between the obtained categories and the ground-truth categories.
@@ 5.2 Baseline Methods and Parameter Settings
Deep Learning-based Methods. (a) CommDGI learns node embeddings by maximizing the mutual information between local node representations and global graph summaries. (b) DGCluster transforms modularity maximization into a differentiable loss function, which is jointly optimized with GNN-based representation learning in an end-to-end manner. (c) DCGL adopts a pseudo-siamese network architecture, extracting features from both structural and attribute perspectives, and enhances representations through cross-perspective contrastive learning. (d) MAGI uses the modularity matrix as an anchor point for contrastive learning, capturing higher-order structural similarities without the need for graph augmentation. (e) GCN designs multi-hop graph convolutions to adaptively fuse high-order neighborhood information, learning more comprehensive node representations. (f) CPGCL jointly learns node representations and community soft assignments, dynamically optimizing sample pairs in contrastive learning to mitigate the issue of false negatives.
@@
The framework consists of the GCN-CDM and TSMM, with the number of training epochs set to 20. The GCN-CDM employs a single-layer convolutional network for node representation learning. The hyperparameter $\delta$ is set to 30. The Adam optimizer is used for 300 iterations of model training, with a learning rate of 0.001 and a weight decay of 0.005. For the TSMM, the AdamW optimizer is used, with the following learning rates for different datasets: 5e-5 for the Cora dataset, 5e-6 for Citeseer, WikiCS, and Instagram datasets, and 5e-7 for Pubmed and Arxiv_2023 datasets. A cosine learning rate scheduler with a warmup ratio of 0.001 is employed to smooth the start and facilitate annealing convergence. The number of communities in all experiments is set to the number of communities detected in the experiments conducted in this paper.
The parameters for the deep learning-based comparative experiments are set according to the original papers to ensure algorithm performance. For example, the learning rate for CommDGI is set to 0.001 with 500 iterations. The learning rate for DGCluster is set to 0.001 with 300 iterations. The learning rate for DCGL is set to 0.001 with 300 iterations. The learning rate for MAGI is set to 0.0005 with 400 iterations. The learning rate for MGCN is set to 0.003 with 700 iterations. The learning rate for CPGCL is set to 0.0007 with 600 iterations.
@@ 5.3 Experimental Results
Table 1: Performance comparison of different algorithms. ’OM’ indicates an out of memory error, ’N/A’ indicates the algorithm runtime exceeds five days, and ’NAN’ indicates the algorithm encountered a NAN error.
min DBI: max Q: max NMI: max ACC: max F1: max ARI:
@@ 5.4 MCDM Ranking
To comprehensively evaluate the overall performance of different algorithms, this study employs the TOPSIS multi-criteria decision-making (MCDM) method. Given the diversity of the six evaluation metrics selected in this experiment and the strong correlation between some of the metrics, the CRITIC method is used to automatically calculate objective weights, thereby avoiding subjective bias introduced by manual weighting. The specific process is as follows: First, the raw evaluation metrics are processed with special markings (e.g., OM, N/A and NAN) and normalized to eliminate differences in dimensions and directions. Then, based on the CRITIC method, weights are determined by considering the metrics’ distinguishability and information independence. Finally, the TOPSIS method is applied to calculate the proximity coefficients of each algorithm to the positive ideal solution and the negative ideal solution. Algorithms are ranked based on these coefficients, providing an objective and unified performance ranking under the multi-dimensional evaluation system. The results are shown in Table 2.
The experimental results show that, on multiple real-world datasets, the proposed method consistently reflects the overall performance differences between algorithms. The method presented in this paper ranks first in most datasets, validating its advantage in multi-metric comprehensive performance.
@@ 5.5 Analysis of Mamba Classification Performance Driven by Community Detection
max ACC: max F1: max ARI: Cora Community Pseudo-label Supervision 0.768095 0.782111 0.584886 10% Real Label Supervision Training 0.318685 0.127400 0.069188 30% Real Label Supervision Training 0.266617 0.142171 0.031145 50% Real Label Supervision Training 0.317947 0.108567 0.062094 Citeseer Community Pseudo-label Supervision 0.677966 0.638612 0.450242 10% Real Label Supervision Training 0.242624 0.111425 0.074428 30% Real Label Supervision Training 0.278406 0.117702 0.087771 50% Real Label Supervision Training 0.279033 0.087257 0.104749 Wikics Community Pseudo-label Supervision 0.615588 0.603117 0.502681 10% Real Label Supervision Training 0.249039 0.069054 0.103249 30% Real Label Supervision Training 0.496111 0.010597 0.354123 50% Real Label Supervision Training 0.603531 0.059397 0.507738 Pubmed Community Pseudo-label Supervision 0.678957 0.679474 0.295725 10% Real Label Supervision Training 0.718162 0.292692 0.425924 30% Real Label Supervision Training 0.871735 0.228584 0.735780 50% Real Label Supervision Training 0.877923 0.206319 0.759428 Arixv_2023 Community Pseudo-label Supervision 0.516278 0.417535 0.402997 10% Real Label Supervision Training 0.338478 0.020023 0.241136 30% Real Label Supervision Training 0.507684 0.011970 0.422863 50% Real Label Supervision Training 0.500134 0.009026 0.531469 Instagram Community Pseudo-label Supervision 0.441838 0.367109 0.010666 10% Real Label Supervision Training 0.110768 0.047447 0.002642 30% Real Label Supervision Training 0.105212 0.066849 0.002381 50% Real Label Supervision Training 0.129729 0.059000 0.003884
@@ 5.8.1 Ablation Analysis of TSMM
min DBI: max Q: max NMI: max ACC: max F1: max ARI: Cora Without TSLL 0.518072 0.729730 0.567524 0.740030 0.755850 0.534452 With TSLL 0.162303 0.725636 0.603979 0.768095 0.782111 0.584886 Citeseer Without TSLL 0.432800 0.777627 0.434697 0.612053 0.580937 0.432320 With TSLL 0.324130 0.723163 0.432592 0.677966 0.638612 0.450242 Wikics Without TSLL 0.710040 0.628633 0.503028 0.604307 0.585585 0.479505 With TSLL 0.415174 0.628361 0.512972 0.615588 0.603117 0.502681 Pubmed Without TSLL 0.581284 0.650354 0.293325 0.619973 0.610863 0.266285 With TSLL 0.227610 0.563854 0.318398 0.678957 0.679474 0.295725 Arxiv_2023 Without TSLL 1.095555 0.460959 0.204377 0.351162 0.246508 0.198183 With TSLL 0.567635 0.530952 0.350075 0.516278 0.417535 0.402997 Instagram Without TSLL 0.684267 0.513729 0.007804 0.416703 0.333898 0.014957 With TSLL 0.575601 0.470651 0.006813 0.441838 0.367109 0.010666
@@ 5.8.2 Ablation Analysis of GCN-CDM
min DBI: max Q: max NMI: max ACC: max F1: max ARI: Cora Without GCN-CDM 0.200989 0.723327 0.598183 0.761448 0.773862 0.585462 With GCN-CDM 0.162303 0.725636 0.603979 0.768095 0.782111 0.584886 Citeseer Without GCN-CDM 1.440825 0.55883 0.370513 0.553045 0.515063 0.366366 With GCN-CDM 0.32413 0.723163 0.432592 0.677966 0.638612 0.450242 Wikics Without GCN-CDM 0.915159 0.596449 0.496284 0.60106 0.575458 0.476269 With GCN-CDM 0.415174 0.628361 0.512972 0.615588 0.603117 0.502681 Pubmed Without GCN-CDM 0.956297 0.519531 0.313856 0.66856 0.606519 0.293834 With GCN-CDM 0.22761 0.563854 0.318398 0.678957 0.679474 0.295725 Arxiv_2023 Without GCN-CDM 1.666361 0.277522 0.311058 0.349496 0.319309 0.247112 With GCN-CDM 0.567635 0.530952 0.350075 0.516278 0.417535 0.402997 Instagram Without GCN-CDM 1.65179 0.212626 0.004092 0.182997 0.232552 0.000654 With GCN-CDM 0.575601 0.470651 0.006813 0.441838 0.367109 0.010666

### v2

@@ Abstract
Pretrained language models offer strong text understanding capabilities but remain difficult to deploy in real-world text-attributed networks due to their heavy dependence on labeled data. Meanwhile, community detection methods typically ignore textual semantics, limiting their usefulness in downstream applications such as content organization, recommendation, and risk monitoring. To overcome these limitations, we present Dual Refinement Cycle Learning (DRCL), a fully unsupervised framework designed for practical scenarios where no labels or category definitions are available. DRCL integrates structural and semantic information through a warm-start initialization and a bidirectional refinement cycle between a GCN-based Community Detection Module (GCN-CDM) and a Text Semantic Modeling Module (TSMM). The two modules iteratively exchange pseudo-labels, allowing semantic cues to enhance structural clustering and structural patterns to guide text representation learning without manual supervision. Across several text-attributed graph datasets, DRCL consistently improves the structural and semantic quality of discovered communities. Moreover, a Mamba-based classifier trained solely from DRCL’s community signals achieves accuracy comparable to supervised models, demonstrating its potential for deployment in large-scale systems where labeled data are scarce or costly. The code is available at https://github.com/wuanghoong/DRCL.git.
@@
[a] organization=College of Physics and Information Engineering, Minnan Normal University, city=Zhangzhou, state=Fujian, postcode=363000, country=China
@@ 3 Preliminaries
Definition 1. Text-Attributed Graphs. A text-attributed graph is represented as a quadruple $G=(V,E,X,T)$. Here $V=\{v_{1},v_{2},...,v_{n}\}$ is the node set and $|V|=n$ is the number of nodes. The edge set $E\subseteq V\times V$ describes pairwise relations: an edge $e_{ij}=(v_{i},v_{j})\in E$ indicates that nodes $v_{i}$ and node $v_{j}$ are adjacent, and we denote the total number of edges by $|E|=M$. The matrix $X=\left\{x_{1},x_{2},x_{3},\ldots,x_{n}\right\}\in\mathbb{R}^{n\times d}$ collects the attribute features of all nodes, where the $i$-th row vector $x_{i}\in\mathbb{R}^{d}$ is the $d$-dimensional feature representation of node $v_{i}$. The set $T=\{t_{1},t_{2},...,t_{n}\}$ stores the textual content attached to the nodes, and each $t_{i}$ corresponds to the text associated with node $v_{i}$. The combinatorial structure of $G$ is encoded by the adjacency matrix $A\in\{0,1\}^{n\times n}$ with entries $a_{ij}=1$ if $(v_{i},v_{j})\in E$ and $a_{ij}=0$ otherwise. From $A$ we obtain the diagonal degree matrix $D=\operatorname{diag}\left(d_{1},d_{2},d_{3},\ldots,d_{n}\right)$, where each diagonal element $d_{i}=\sum_{j=1}^{n}a_{ij}$ counts the number of neighbors directly connected to node $v_{i}$.
Definition 2. Graph Convolutional Networks. Graph Convolutional Networks (GCNs) implement neural message passing on graphs and can be viewed as a particular form of spectral filtering of graph signals.
The commonly used propagation rule arises from a first-order Chebyshev polynomial approximation of the spectral graph convolution, leading to an efficient neighborhood aggregation scheme in the spatial domain. The update rule of GCN is
@@
where $H^{(l)}=\{h_{1}^{(l)},h_{2}^{(l)},...,h_{n}^{(l)}\}\in\mathbb{R}^{n\times d^{\prime}}$ denotes the node representation matrix of the $l$-layer, and the input of the first layer is $H^{(0)}=X$. The adjacency matrix with self-loops is $\hat{A}=A+I$, where $I$ is the identity matrix, ensuring that each node also propagates its own feature during aggregation. The corresponding degree matrix $\hat{D}$ is diagonal with entries $\hat{D}_{ii}={\textstyle\sum_{j}}\hat{A}_{ij}$. The normalized adjacency matrix is then defined as $\tilde{A}=\hat{D}^{-1/2}\hat{A}\hat{D}^{-1/2}$ which stabilizes the training process and alleviates gradient explosion or vanishing. The matrix $W^{(l)}$ is the learnable weight parameter of the $l$-th layer, and $\sigma(\cdot)$ denotes a nonlinear activation function (PReLU in our implementation), enhancing the model’s expressive power and the discriminability of node representations.
@@ 5.1 Datasets and Evaluation Metrics
To ensure the completeness and comparability of the evaluation, this paper selects six widely used metrics for comprehensive assessment: NMI, ACC, F1, ARI, DBI and Q. Among these, DBI and Q measure the quality of community partitioning from the perspectives of internal clustering cohesion and graph structure consistency, respectively. NMI, ACC, F1, and ARI are external consistency metrics commonly used in both community detection and language text classification tasks, reflecting the degree of alignment between the obtained categories and the ground-truth categories.
@@ 5.2 Baseline Methods and Parameter Settings
Deep Learning-based Methods. (a) In CommDGI, node embeddings are learned by employing a mutual-information objective between local node-level representations and a global summary vector of the whole graph. (b) DGCluster casts modularity maximization as a differentiable loss term, which is optimized together with a GNN encoder so that representation learning and clustering are trained in a single end-to-end pipeline. (c) DCGL builds a pseudo-siamese network with two branches that separately encode structural information and attribute information; the two views are then aligned by cross-view contrastive learning to strengthen the final embeddings. (d) MAGI regards the modularity matrix as an anchor in a contrastive learning framework, enabling the model to capture high-order structural similarity without explicitly constructing augmented graphs. (e) MGCN introduces multi-hop graph convolution modules that adaptively aggregate signals from higher-order neighborhoods, leading to richer and more informative node representations. (f) CPGCL jointly optimizes node embeddings and soft community assignment variables, and continuously reselects contrastive sample pairs during training so as to alleviate the problem of false negatives.
@@
The proposed framework is composed of GCN-CDM and TSMM and is trained for 20 epochs. GCN-CDM uses a single-layer graph convolution network to learn node representations, with the hyperparameter $\delta$ fixed at 30. Model parameters are optimized with Adam for 300 update steps, using a learning rate of 0.001 and a weight decay of 0.005. For TSMM, we adopt the AdamW optimizer and choose dataset-specific learning rates: 5e-5 for the Cora dataset, 5e-6 for Citeseer, WikiCS, and Instagram datasets, and 5e-7 for Pubmed and Arxiv_2023 datasets. A cosine learning-rate schedule with a warmup ratio of 0.001 is applied to stabilize the initial phase and to enable smooth annealing. In all experiments, the number of communities is set equal to the number detected by our framework on each dataset.
For deep learning–based baselines, we closely follow the hyperparameter settings reported in the original works to ensure fair comparison. Specifically, CommDGI is trained with a learning rate of 0.001 for 500 iterations; DGCluster and DCGL both use a learning rate of 0.001 with 300 iterations; MAGI adopts a learning rate of 0.0005 with 400 iterations; MGCN uses a learning rate of 0.003 with 700 iterations; and CPGCL is trained with a learning rate of 0.0007 for 600 iterations.
@@ 5.3 Experimental Results
[…]
DBI($\downarrow$) Q($\uparrow$) NMI($\uparrow$) ACC($\uparrow$) F1($\uparrow$) ARI($\uparrow$)
@@ 5.4 MCDM Ranking
To assess the algorithms from a holistic perspective, we adopt the TOPSIS-based multi-criteria decision-making (MCDM) framework. Since the six evaluation indicators used in our experiments differ in type and exhibit non-negligible correlations, we employ the CRITIC method to automatically derive objective weights, thereby avoiding subjective bias from manual assignment. The procedure is as follows. First, the raw metric values are preprocessed by handling special symbols (e.g., OM, N/A, NAN) and then normalized to remove inconsistencies in scale and optimization direction. Next, CRITIC is applied to compute the weight of each metric by jointly considering its discriminative power and information redundancy. Finally, the TOPSIS scheme is used to compute the closeness of each algorithm to the positive and negative ideal solutions, and the algorithms are ranked according to these proximity scores, yielding a unified and objective performance ordering under the multi-dimensional evaluation system. The detailed results are reported in Table 2.
Experimental evidence on multiple real-world datasets shows that this evaluation pipeline can reliably capture the overall performance gaps among competing methods. The approach proposed in this study achieves the top rank on most datasets, demonstrating its clear advantage in terms of aggregated performance across multiple metrics.
@@ 5.5 Analysis of Mamba Classification Performance Driven by Community Detection
ACC($\uparrow$) F1($\uparrow$) ARI($\uparrow$) Cora Community Pseudo-label Supervision 0.768095 0.782111 0.584886 10% Real Label Supervision Training 0.318685 0.127400 0.069188 30% Real Label Supervision Training 0.266617 0.142171 0.031145 50% Real Label Supervision Training 0.317947 0.108567 0.062094 Citeseer Community Pseudo-label Supervision 0.677966 0.638612 0.450242 10% Real Label Supervision Training 0.242624 0.111425 0.074428 30% Real Label Supervision Training 0.278406 0.117702 0.087771 50% Real Label Supervision Training 0.279033 0.087257 0.104749 Wikics Community Pseudo-label Supervision 0.615588 0.603117 0.502681 10% Real Label Supervision Training 0.249039 0.069054 0.103249 30% Real Label Supervision Training 0.496111 0.010597 0.354123 50% Real Label Supervision Training 0.603531 0.059397 0.507738 Pubmed Community Pseudo-label Supervision 0.678957 0.679474 0.295725 10% Real Label Supervision Training 0.718162 0.292692 0.425924 30% Real Label Supervision Training 0.871735 0.228584 0.735780 50% Real Label Supervision Training 0.877923 0.206319 0.759428 Arixv_2023 Community Pseudo-label Supervision 0.516278 0.417535 0.402997 10% Real Label Supervision Training 0.338478 0.020023 0.241136 30% Real Label Supervision Training 0.507684 0.011970 0.422863 50% Real Label Supervision Training 0.500134 0.009026 0.531469 Instagram Community Pseudo-label Supervision 0.441838 0.367109 0.010666 10% Real Label Supervision Training 0.110768 0.047447 0.002642 30% Real Label Supervision Training 0.105212 0.066849 0.002381 50% Real Label Supervision Training 0.129729 0.059000 0.003884
@@ 5.8.1 Ablation Analysis of TSMM
DBI($\downarrow$) Q($\uparrow$) NMI($\uparrow$) ACC($\uparrow$) F1($\uparrow$) ARI($\uparrow$) Cora Without TSLL 0.518072 0.729730 0.567524 0.740030 0.755850 0.534452 With TSLL 0.162303 0.725636 0.603979 0.768095 0.782111 0.584886 Citeseer Without TSLL 0.432800 0.777627 0.434697 0.612053 0.580937 0.432320 With TSLL 0.324130 0.723163 0.432592 0.677966 0.638612 0.450242 Wikics Without TSLL 0.710040 0.628633 0.503028 0.604307 0.585585 0.479505 With TSLL 0.415174 0.628361 0.512972 0.615588 0.603117 0.502681 Pubmed Without TSLL 0.581284 0.650354 0.293325 0.619973 0.610863 0.266285 With TSLL 0.227610 0.563854 0.318398 0.678957 0.679474 0.295725 Arxiv_2023 Without TSLL 1.095555 0.460959 0.204377 0.351162 0.246508 0.198183 With TSLL 0.567635 0.530952 0.350075 0.516278 0.417535 0.402997 Instagram Without TSLL 0.684267 0.513729 0.007804 0.416703 0.333898 0.014957 With TSLL 0.575601 0.470651 0.006813 0.441838 0.367109 0.010666
@@ 5.8.2 Ablation Analysis of GCN-CDM
DBI($\downarrow$) Q($\uparrow$) NMI($\uparrow$) ACC($\uparrow$) F1($\uparrow$) ARI($\uparrow$) Cora Without GCN-CDM 0.200989 0.723327 0.598183 0.761448 0.773862 0.585462 With GCN-CDM 0.162303 0.725636 0.603979 0.768095 0.782111 0.584886 Citeseer Without GCN-CDM 1.440825 0.55883 0.370513 0.553045 0.515063 0.366366 With GCN-CDM 0.32413 0.723163 0.432592 0.677966 0.638612 0.450242 Wikics Without GCN-CDM 0.915159 0.596449 0.496284 0.60106 0.575458 0.476269 With GCN-CDM 0.415174 0.628361 0.512972 0.615588 0.603117 0.502681 Pubmed Without GCN-CDM 0.956297 0.519531 0.313856 0.66856 0.606519 0.293834 With GCN-CDM 0.22761 0.563854 0.318398 0.678957 0.679474 0.295725 Arxiv_2023 Without GCN-CDM 1.666361 0.277522 0.311058 0.349496 0.319309 0.247112 With GCN-CDM 0.567635 0.530952 0.350075 0.516278 0.417535 0.402997 Instagram Without GCN-CDM 1.65179 0.212626 0.004092 0.182997 0.232552 0.000654 With GCN-CDM 0.575601 0.470651 0.006813 0.441838 0.367109 0.010666